OpenStudy (anonymous):

s the anyone know how can i rewrite this f(x)=x^2−2x+3 in the form of f(x)=a(x-h)^2+k

OpenStudy (phi):

you "complete the square" using the idea that (x + g)(x+g)= x^2 +2g x + g^2 match your expression x^2 +2g x + g^2 x^2 -2 x + 3 specifically, match the 2nd terms: 2g x = -2 x you find g= -1 and g^2 = 1 write x^2 - 2 x + 1 - 1 + 3 I added and subtracted 1 (x^2-2x + 1) + 3-1 (x-1)^2 + 2